Change the values of acceleration in Regions A, B, C, D, and the initial Velocity so that the person reaches the 40 m mark in 10 seconds. Play the simulation so you know your answers wil be correct

What are your values for (5 answers)

Acceleration in A,

Acceleration in B,

Acceleration in C,

Acceleration in D,

and Initial Velocity
Answer #1

Assuming that the person is moving with initial velocity 0.

Let A, B, C and D are four points on the line or the track.

A) V= displacement/ time taken

=10/10

=1 m/s.

a .=1/10

= 0.1 m/s2

B) V= 20/10

= 2m/ s

a=2/10

=0.2m/s2

C) V =30/10

. =3m/ s

. a . =3/10

. =0.3m/s2

D) V=40/10

. V=4m/s

. a=4/10

=0.4m/s2

Initial velocity is U = 0.